OpenStudy (chunkymonkey):

Matt 6 – 4(3 – 5)2 + 30 ÷ 5 6 – 4(–2)2 + 30 ÷ 5 6 – 4(4) + 30 ÷ 5 6 – 16 + 30 ÷ 5 −10 + 30 ÷ 5 20 ÷ 5 4 Karen 6 – 4(3 – 5)2 + 30 ÷ 5 6 – 4(–2)2 + 30 ÷ 5 6 – 4(−4) + 30 ÷ 5 6 + 16 + 30 ÷ 5 6 + 16 + 6 22 + 6 28

OpenStudy (chunkymonkey):

whats wrong with them?need help!

OpenStudy (anonymous):

for Matt, do you see what is wrong in this step? 6 – 4(–2)2 + 30 ÷ 5 = 6 – 4(4) + 30 ÷ 5

OpenStudy (anonymous):

Ok in matt's I already see a problem in step 3 (-2)2 should be -4

OpenStudy (anonymous):

for Karen, do you see how her division is incorrect on this step compared to when matt did his division? 6 + 16 + 30 ÷ 5 = 6 + 16 + 6

OpenStudy (chunkymonkey):

ooohhhh ok i see it thank you!